OpenStudy (anonymous):

An airplane is flying with a velocity of 95.0m/s at an angle of 22.0 degreesabove the horizontal. When the plane is a distance 116m m directly above a dog that is standing on level ground, a suitcase drops out of the luggage compartment. How far from the dog will the suitcase land? ignore air resistance

OpenStudy (anonymous):

Ok, let me see : First you need to decompose the velocity : \( \Large V = 95\) m/s \(\Large V_x = 95\times cos(22) = 88.08\) m/s \(\Large V_y = 95\times sin(22) = 35.59\) m/s The suitcase first will go up, and then it will go down, the suitcase is \(\Large 116\) over the ground, now, let's work with the vertical velocity, let's find the maximum altitude, the time the suitcase takes, and then, let's find the total flying time : \(\Large V_y = 35.59m/s \) maximum altitude : \(\Large Vf^2 = V_y^2 - 2\times g \times h\) considering \(\Large g = 9.8 m/s^2\) When the suitcase reaches its maximum altitude, \(\Large Vf = 0 m/s\) \(\Large 0^2 = 35.59^2 - 2\times9.8\times h\) \(\Large h = 64.62\) meters \(\Large Vf = V_y - g\times t\) \(\Large 0 = 35.59 - 9.8\times t\) \(\Large t = 3.63\) seconds Now, let's find the time, the suitcase took to go down : total altitude \(\Large = 116 + 64.62 = 180.62\) meters \(\Large 180.62 = 1/2\times9.8\times t'^2\) \(\Large t' = 6.07\) seconds Total flying time \(\Large = T = 6.07 + 3.63 = 9.7\) seconds Now, we can find how far the suitcase landed : \(\Large X = V_x\times T\) \(\Large X = 88.08\times9.7 = 854.376\) meters Hope that helps

OpenStudy (anonymous):

thank you so much! @junghyunran